OpenStudy (anonymous):

In 3 minutes, a conveyor belt moves 400 lbs of recyclable aluminum from the delivery truck to a storage area. A smaller belt moves the same quantity of cans the same distance in 4 minutes. If both belts are used, find how long it takes to move the cans to the storage area. I think it is 3 1/2 minutes but I'm not sure. Please help!

OpenStudy (anonymous):

@shamim

OpenStudy (anonymous):

@angel12310 @fruit.snacks @EclipsedStar @ILUVSOCCER @Deadpool117 @marylou004 @Quan99 @gottennis121 @wolfd.andi @e.mccormick @Javk @pitamar @pootypie

OpenStudy (anonymous):

you to connections

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@Agl202

OpenStudy (anonymous):

\[\frac{400}{3}+\frac{400}{4}=\frac{800}{x} \]x = 24/7=3.42857 Been a while since I addressed this type of rate problem.

OpenStudy (anonymous):

wait, wouldn't first find the common denominator which is 12x? than multiply that by those three fractions?

OpenStudy (anonymous):

nvm i got it

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Thank you for the medal.